### v1

@@ Abstract
Scientific Workflow Systems (SWSs) are advanced software frameworks that drive modern research by orchestrating complex computational tasks and managing extensive data pipelines. These systems offer a range of essential features, including modularity, abstraction, customization, interoperability, workflow composition tools, resource management, error handling, and comprehensive documentation. Utilizing these frameworks accelerates the development of scientific computing, resulting in more efficient and reproducible research outcomes. Despite their significance, developing a user-friendly, efficient, and adaptable SWS poses several challenges that are not always well-documented or understood. This study explores these challenges through an in-depth analysis of interactions on Stack Overflow (SO) and GitHub, key platforms where developers and researchers discuss and resolve issues. In particular, we leverage topic modeling (BERTopic) to understand the topics SWSs developers discuss on these platforms. Then, we examine the popularity and difficulty of those topics. We identified 10 topics developers discuss on SO (e.g., Workflow Creation and Scheduling, Data Structures and Operations, Workflow Execution) and found that workflow execution is the most challenging among them. By analyzing GitHub issues, we identified 13 topics (e.g., Errors and Bug Fixing, Documentation, Dependencies) and discovered that data structures and operations is the most difficult topics in this context. We also found common topics between SO and GitHub, such as data structures and operations, task management, and workflow scheduling. Additionally, we categorized each topic by type (How, Why, What, and Others) and observed that the How type consistently dominates across all topics, indicating a need for procedural guidance among developers. This dominance of the How type is also prevalent in other domains, such as Chatbots and Mobile development. We believe that our study will guide future research in proposing tools and techniques to help the community overcome the challenges developers face when developing SWSs.
@@ 1 Introduction
Scientific Workflow Systems (SWSs) are software frameworks designed to streamline the design, execution, and management of complex, data-intensive scientific experiments and research [1]. They allow users to create, visualize, and automate workflows composed of interconnected steps, each carrying out specific data operations. These systems orchestrate tasks, manage data dependencies, and ensure reproducibility by consistently capturing and executing workflows [2]. By enhancing the efficiency of running and analyzing experiments, they facilitate collaboration and meet the rigorous demands of scientific research. SWSs serve as the bedrock upon which efficient, transparent, and impactful research is built, propelling knowledge and innovation across diverse fields. They accomplish mundane yet essential tasks intrinsic to scientific inquiry—ranging from data acquisition to analysis and reporting. By liberating researchers from time-consuming tasks, SWSs allow them to focus on more complex and creative work, which helps speed up and improve research results [3]. They provide a platform to create and distribute the necessary computational steps for tasks involving data analysis and simulations, all while concealing the intricate infrastructural complexities [4]. More importantly, they capture the entire scientific experimentation process, offering valuable insights for reproducing, reusing, or adapting these procedures.
@@
Traditionally, various types of scientific analyses are implemented by practitioners using either scripting languages (Python, Perl, or Bash) requiring fairly sophisticated computational skills or by manually calling and storing intermediate results, which requires a considerable amount of time organizing data files and managing the overall analysis process [5]. Alternatively, SWSs can help practitioners by providing higher-level modeling approaches for explicitly specifying analyses as well as by providing several generic services for optimizing workflow execution, visualizing workflows and workflow results, managing intermediate data products, and storing the details of past workflow executed results [6]. Their ability to handle large volumes of data and integrate various computational tools makes them essential for advancing research across diverse scientific domains. Furthermore, SWSs afford researchers, developers, and project managers real-time insights into the progress of research endeavors, empowering them to identify bottlenecks, allocate resources judiciously, and optimize workflow execution. The customizable nature of SWSs allows practitioners to tailor workflows to their specific needs and preferences, enhancing their utility and versatility. In essence, SWSs emerge not merely as tools of convenience but as indispensable allies in the relentless pursuit of scientific excellence.
Figure 1: Year-wise Number of Posts, Issues, and Pull Requests
@@
Practitioners encountered a significant deficiency in usability, effectiveness, and reliability when dealing with these SWSs [15]. Furthermore, the consistent rise in the number of posts, issues, and pull requests (for selected SWSs) shown in Figure 1 implies a growing demand for increased support for developers engaged in SWSs. Given the widespread use of SWSs across various domains and the interest they have garnered, including from the software analytics community, it becomes crucial to comprehend these challenges for advancing the efficiency and effectiveness of these systems, which are integral to data-intensive research and complex scientific computations. Practitioners often encounter unique difficulties in SWSs, such as managing heterogeneous data sources, ensuring reproducibility, and optimizing workflows for high-performance computing environments. By systematically identifying these challenges, researchers and practitioners can tailor tools and methodologies to better support the development of robust, scalable, and user-friendly workflows. Addressing these challenges not only enhances the usability and reliability of SWSs but also accelerates scientific discovery by enabling researchers to focus on their domain-specific tasks rather than the intricacies of the workflow system itself. Moreover, understanding these challenges provides valuable insights that can drive innovation in SWSs design, leading to more efficient computational solutions that meet the evolving needs of the scientific community [7]. Therefore, we have analyzed SO posts, GitHub issues, and pull requests related to SWSs to identify developer challenges in our research.
@@
In this paper, we conduct an empirical investigation to rigorously understand the needs and desiderata of developers working with SWSs. Our objective is to gain a precise understanding of the challenges developers face and identify potential areas for improvement, aiming to pinpoint specific issues and obstacles encountered in their work. To achieve this, we have compiled a dataset of 35,619 SO posts and 140,400 GitHub issues and pull requests related to SWSs. Using a combination of manual analysis and automated topic modeling, we examine these SO posts and GitHub data to uncover the specific challenges developers encounter. In particular, our study aims to answer the following three research questions (RQs):
@@
  • RQ1: What topics are SWSs developers asking about on Stack Overflow, and which topics are raised in the GitHub projects? The SWSs-related posts, issues, and pull requests may reflect developers’ challenges when learning or developing an SWS. To understand these challenges, we utilize topic modeling to extract semantic topics from SWSs-related data on SO and GitHub. From SO posts, we identified 10 topics, which include traditional software engineering topics (e.g., data structures and operations, dependencies management) and SWSs-specific themes (e.g., workflow creation and scheduling, workflow execution). Similarly, analysis of GitHub data yielded 13 topics, encompassing traditional software engineering topics (e.g., errors and bug fixing, dependencies) and SWSs-specific topics (e.g., data management and task scheduling, annotations). By comparing the challenges developers face on SO and GitHub, we aim to identify any similarities. By answering RQ1, we seek to pinpoint the areas in SWSs development where developers frequently seek assistance, thereby highlighting the aspects that commonly pose challenges for them.
@@
  • RQ2: What types of questions are SWSs developers asking? Developers ask different (i.e., How, Why) types of questions to address different challenges. To answer RQ2, we adopt a methodology similar to previous research on identifying types of posts on Stack Overflow [16, 17, 18]. We select a statistically significant number of posts for each topic and classify them into types such as How, What, Why, and Others. This approach enables us to systematically analyze the nature of the questions developers ask and gain deeper insights into their needs and challenges. Our analysis shows that SWSs developers primarily seek guidance on specific implementation routines, workflow examples, and troubleshooting—highlighting the dominance of How questions, similar to other domains like ChatBot and Mobile Development. Our findings underscore the need for step-by-step instructions and real-world scenarios to help developers achieve their goals in SWSs implementation.
@@
  • RQ3: To what extent do developers perceive the revealed challenges in terms of difficulties? To assess the difficulty of each topic, we use two metrics: the percentage of unresolved posts/issues/pull requests and the median time to resolve them. This approach follows the methodologies of prior studies [16, 19, 20, 21, 22, 17]. In RQ3, our goal is to identify the most challenging topics to resolve, as these are likely to pose significant difficulties for all but the most experienced developers. We identify workflow execution as the most challenging topic on SO due to its complexity in managing distributed resources, handling large-scale data processing, ensuring fault tolerance, and optimizing performance across diverse environments. Similarly, on GitHub, Kubernetes deployments for SWSs frameworks pose significant challenges, primarily due to difficulties in ensuring scalability, fault tolerance, and efficient resource allocation required for managing complex and distributed workflows. This analysis highlights areas within SWSs development where additional support and resources are essential.
@@
Paper Organization: The remainder of this paper is structured as outlined as follows: In Section 2, we briefly discuss Scientific Workflows, Scientific Workflow Systems(SWSs), and related work. Moving on to Section 3, we detail our study’s methodology. Our findings are presented in Section 4, and Section 5 illustrates the implications. Section 6 addresses potential threats to the validity of our results. Finally, Section 7 concludes the paper, highlighting directions for future research.
@@ 2 Background and Related Work
SWSs emerged as a response to the growing complexity and scale of scientific research [23]. As computational power increased and data generation exploded across various scientific domains, researchers needed more efficient ways to manage, automate, and reproduce their experiments. Traditional manual methods were inadequate for handling the intricate sequences of data processing tasks required. SWSs provide a structured framework for designing, executing, and sharing multi-step computational processes. They integrate diverse tools and data sources, facilitating collaboration and enhancing the reproducibility of scientific research. We aim to understand developers’ challenges in developing SWSs by analyzing SO posts, GitHub issues, and pull requests. In this section, we provide an overview of the background and previous research associated with SWSs development. Additionally, we explore prior work that utilized topic modeling techniques to gain insights into developer perspectives.
@@ 2.2 Prior Studies on SWSs
Numerous studies have been conducted to improve the usability, reliability, and scalability of SWSs. For example, Molder et al. [25] focus on sustainable data analysis with Snakemake. Reproducibility challenges are explored in [43], while [1] discourse reusability issues. Collaborative enhancements are discussed in [44], and Jagla et al. [45] extend SWSs for next-generation sequencing analysis. Evaluation criteria for SWSs are examined in [46, 47]. Additionally, Di Tommaso et al. address reliability in Nextflow [34], and studies on common workflow language (CWL) are discussed in [40, 48]. Jain et al. discuss SWSs for high-throughput applications in [49], while Ahmad et al. explore workflow scheduling challenges in [50]. Hundreds of SWSs have been developed to aid scientific workflow operations [51], and extensive research ([52, 53, 54, 55, 15]) has been conducted to identify the challenges associated with the efficient construction and utilization of SWSs. Several studies [56, 57, 58] have explored the intersection of SWSs and large language models (LLMs). Integrating LLMs, such as GPT models, into SWSs can drive substantial advancements across various domains, particularly in automating workflow design, optimizing data-driven processes, and enhancing scientific collaboration. This cutting-edge AI and SWSs combination could lead to more efficient, accessible, and scalable workflows. Unlike previous studies, this work makes the first attempt to understand the challenges of SWSs as perceived by developers.

### v2

@@ Abstract
Scientific Workflow Systems (SWSs) are advanced software frameworks that drive modern research by orchestrating complex computational tasks and managing extensive data pipelines. These systems offer a range of essential features, including modularity, abstraction, customization, interoperability, workflow composition tools, resource management, error handling, and comprehensive documentation. Utilizing these frameworks accelerates the development of scientific computing, resulting in more efficient and reproducible research outcomes. Despite their significance, developing a user-friendly, efficient, and adaptable SWS poses several challenges that are not always well-documented or understood. This study explores these challenges through an in-depth analysis of interactions on Stack Overflow (SO) and GitHub, key platforms where developers and researchers discuss and resolve issues. In particular, we leveraged topic modeling (BERTopic) to understand the topics SWSs developers discuss on these platforms. Then, we examined the popularity and difficulty of those topics. We identified 10 topics developers discuss on SO (e.g., Workflow Creation and Scheduling, Data Structures and Operations, Workflow Execution) and found that workflow execution is the most challenging among them. By analyzing GitHub issues, we identified 13 topics (e.g., Errors and Bug Fixing, Documentation, Dependencies) and discovered that errors and bug fixing is the most dominant topics in this context. We found system redesign and API migration to be the most challenging topics utilizing GitHub data. A cross-platform comparison revealed overlapping concerns such as task management and data operations. Additionally, we categorized each topic by type (How, Why, What, and Others) and observed that the How type consistently dominates across all topics, indicating a need for procedural guidance among developers. This dominance of the How type is also prevalent in other domains, such as Chatbots and Mobile development. We believe that our study will guide future research in proposing tools and techniques to help the community overcome the challenges developers face when developing SWSs.
@@ 1 Introduction
Scientific Workflow Systems (SWSs) are software frameworks designed to streamline the design, execution, and management of complex, data-intensive scientific experiments and research [1]. They allow users to create, visualize, and automate workflows composed of interconnected steps, each carrying out specific data operations. These systems orchestrate tasks, manage data dependencies, and ensure reproducibility by consistently capturing and executing workflows [2]. By enhancing the efficiency of running and analyzing experiments, they facilitate collaboration and meet the rigorous demands of scientific research. SWSs serve as the bedrock upon which efficient, transparent, and impactful research is built, propelling knowledge and innovation across diverse fields. They accomplish mundane yet essential tasks intrinsic to scientific inquiry—ranging from data acquisition to analysis and reporting. By liberating researchers from time-consuming tasks, SWSs allow them to focus on more complex and creative work, which helps speed up and improve research results [3]. They provide a platform to create and distribute the necessary computational steps for tasks involving data analysis and simulations, all while concealing the intricate infrastructural complexities [4]. More importantly, they capture the entire scientific experimentation process, offering valuable insights for reproducing, reusing, or adapting these procedures. SWSs such as Galaxy [5], KNIME [6], Snakemake [7], and Nextflow [8] support the design, execution, and management of complex scientific workflows across various domains (e.g., Genomics and Bioinformatics, Microbiology and Virology, Computational Biology, Epidemiology and Public Health). These systems offer diverse interfaces ranging from user-friendly graphical tools (e.g., Galaxy, KNIME) to script- and DSL-based platforms (e.g., Airflow, Snakemake, Nextflow), enabling reproducible, scalable, and efficient computational research.
@@
Traditionally, various types of scientific analyses are implemented by practitioners using either scripting languages (Python, Perl, or Bash) requiring fairly sophisticated computational skills or by manually calling and storing intermediate results, which requires a considerable amount of time organizing data files and managing the overall analysis process [9]. Alternatively, SWSs can help practitioners by providing higher-level modeling approaches for explicitly specifying analyses as well as by providing several generic services for optimizing workflow execution, visualizing workflows and workflow results, managing intermediate data products, and storing the details of past workflow executed results [10]. Lin et al. [11] defined scientific workflow as a computerized facilitation or automation of a scientific process, in whole or part, which usually streamlines a collection of scientific tasks with data channels and dataflow constructs to automate data computation and analysis to enable and accelerate scientific discovery. The ability of SWSs to handle large volumes of data and integrate various computational tools makes them essential for advancing research across diverse scientific domains. Furthermore, SWSs afford researchers, developers, and project managers real-time insights into the progress of research endeavors, empowering them to identify bottlenecks, allocate resources judiciously, and optimize workflow execution. The customizable nature of SWSs allows practitioners to tailor workflows to their specific needs and preferences, enhancing their utility and versatility. In essence, SWSs emerge not merely as tools of convenience but as indispensable allies in the relentless pursuit of scientific excellence.
Figure 1: Year-wise distribution of posts on Stack Overflow, and issues & pull requests on GitHub related to selected SWSs, reflecting the growing developer engagement and activity in SWS development over time.
@@
Practitioners encountered a significant deficiency in usability, effectiveness, and reliability when dealing with these SWSs [20]. Furthermore, the consistent rise in the number of posts, issues, and pull requests (for selected SWSs) shown in Figure 1 implies a growing demand for increased support for developers engaged in SWSs. Given the widespread use of SWSs across various domains and the interest they have garnered, including from the software analytics community, it becomes crucial to comprehend these challenges for advancing the efficiency and effectiveness of these systems, which are integral to data-intensive research and complex scientific computations. Practitioners often encounter unique difficulties in SWSs, such as managing heterogeneous data sources, ensuring reproducibility, and optimizing workflows for high-performance computing environments. By systematically identifying these challenges, researchers and practitioners can tailor tools and methodologies to better support the development of robust, scalable, and user-friendly workflows. Addressing these challenges not only enhances the usability and reliability of SWSs but also accelerates scientific discovery by enabling researchers to focus on their domain-specific tasks rather than the intricacies of the workflow system itself. Moreover, understanding these challenges provides valuable insights that can drive innovation in SWSs design, leading to more efficient computational solutions that meet the evolving needs of the scientific community [12]. Therefore, we analyzed SO posts, GitHub issues, and pull requests related to SWSs to identify developer challenges in our research.
@@
In this paper, we conducted an empirical investigation to rigorously understand the needs and desiderata of developers working with SWSs. Our objective is to gain a precise understanding of the challenges developers face and identify potential areas for improvement, aiming to pinpoint specific issues and obstacles encountered in their work. To achieve this, we compiled a dataset of 35,619 SO posts and 163,118 GitHub issues and pull requests related to SWSs. Using a combination of manual analysis and automated topic modeling, we examined these SO posts and GitHub data to uncover the specific challenges developers encounter. In particular, our study aims to answer the following three research questions (RQs):
@@
  • RQ1: What topics are SWSs developers asking about on Stack Overflow, and which topics are raised in the GitHub projects? The SWSs-related posts, issues, and pull requests may reflect developers’ challenges when learning or developing an SWS. To understand these challenges, we utilized topic modeling to extract semantic topics from SWSs-related data on SO and GitHub. From SO posts, we identified 10 topics, which include traditional software engineering topics (e.g., data structures and operations, dependencies management) and SWSs-specific challenges (e.g., workflow creation and scheduling, workflow execution). Similarly, analysis of GitHub data yielded 13 topics, encompassing traditional software engineering topics (e.g., errors and bug fixing, dependencies) and SWSs-specific topics (e.g.,data shuffling and partitioning strategies, task coordination). By comparing the challenges developers face on SO and GitHub, we aimed to identify any similarities. By answering RQ1, we aimed to pinpoint the areas in SWSs development where developers frequently seek assistance, thereby highlighting the aspects that commonly pose challenges for them.
@@
  • RQ2: What types of questions are SWSs developers asking? Developers ask different (i.e., How, Why) types of questions to address different challenges. To answer RQ2, we adopted a methodology similar to previous research on identifying types of posts on Stack Overflow [21, 22, 23]. We selected a statistically significant number of posts for each topic and classified them into types such as How, What, Why, and Others. This approach enabled us to systematically analyze the nature of the questions developers ask and gain deeper insights into their needs and challenges. Our analysis showed that SWSs developers primarily seek guidance on specific implementation routines, workflow examples, and troubleshooting, highlighting the dominance of How questions, similar to other domains like ChatBot [22] and Mobile Development [21]. Our findings underscored the need for step-by-step instructions and real-world scenarios to help developers achieve their goals in SWSs implementation.
@@
  • RQ3: To what extent do developers perceive the revealed challenges in terms of difficulties? To assess the difficulty of each topic, we used two metrics: the percentage of unresolved posts/issues/pull requests and the median time to resolve them. This approach followed the methodologies of prior studies [21, 24, 25, 26, 27, 22]. In RQ3, our goal is to identify the most challenging topics to resolve, as these are likely to pose significant difficulties for all but the most experienced developers. We identified workflow execution as the most challenging topic on SO due to its complexity in managing distributed resources, handling large-scale data processing, ensuring fault tolerance, and optimizing performance across diverse environments. Similarly, on GitHub, Kubernetes deployments for SWSs frameworks pose significant challenges, primarily due to difficulties in ensuring scalability, fault tolerance, and efficient resource allocation required for managing complex and distributed workflows. This analysis highlighted areas within SWSs development where additional support and resources are essential.
@@
Paper Organization: The remainder of this paper is structured as outlined as follows: In Section 2, we briefly discussed Scientific Workflow Systems(SWSs), Scientific Workflows, Topic Modeling, and related work. Moving on to Section 3, we detailed our study’s methodology. Our findings are presented in Section 4, and Section 5 illustrates the implications. Section 6 addresses potential threats to the validity of our results. Finally, Section 7 concludes the paper, highlighting directions for future research.
@@ 2 Background and Related Work
SWSs emerged as a response to the growing complexity and scale of scientific research [28]. As computational power increased and data generation exploded across various scientific domains, researchers needed more efficient ways to manage, automate, and reproduce their experiments. Traditional manual methods were inadequate for handling the intricate sequences of data processing tasks required. SWSs provide a structured framework for designing, executing, and sharing multi-step computational processes. They integrate diverse tools and data sources, facilitating collaboration and enhancing the reproducibility of scientific research. We aimed to understand developers’ challenges in developing SWSs by analyzing SO posts, GitHub issues, and pull requests. In this section, we provided an overview of the background and previous research associated with SWSs development. Additionally, we explored prior work that utilized topic modeling techniques to gain insights into developer perspectives.
@@ 2.2 Prior Studies on SWSs
SWSs have been widely studied regarding reproducibility, reusability, usability, performance, and collaborative development. These studies provide valuable insights into the capabilities and limitations of current systems, and several findings directly motivate the goals of our work. Several studies highlight that ensuring reproducibility remains one of the most pressing challenges in computational science. For example, reproducibility is often cited as a core benefit of SWSs, yet studies reveal that ensuring reproducibility in practice remains challenging. Molder et al. [29] emphasized the importance of sustainable practices in workflow development, showing that inconsistent environment setups often hamper reproducibility. Studies such as Perez et al. [43] identified that reproducibility issues arise due to poorly documented workflows and unclear execution contexts. Chirigati et al. [44] proposed ReproZip as a solution to improve computational reproducibility, though they acknowledge the lack of standardization across platforms as a limitation.
Reusability is another major concern in SWSs. Garijo et al. [45] proposed the Open Provenance Model for Workflows (OPMW), an end-to-end framework to publish and share workflows using Linked Data principles. Their work showed that many workflows lack rich metadata, making discovery, understanding, and reuse difficult. This finding aligns with our observation that many developer discussions on GitHub and Stack Overflow center on difficulties in reusing or adapting existing workflows due to unclear structure or insufficient context. Similarly, Alam et al. [1] examined the barriers to reusability in scientific workflows, noting that workflow portability across platforms remains a significant challenge.
Collaborative enhancements in SWSs have been explored by the Galaxy project team [46], who emphasized the importance of community-driven infrastructure and tool interoperability to support large-scale, distributed analyses. Jagla et al. [47] demonstrated how SWSs could be extended to handle next-generation sequencing (NGS) tasks, highlighting the need for domain-specific customization and robust data-handling capabilities. Evaluation frameworks for SWSs are reviewed in studies such as Kiran et al. [48] and Sethi et al. [49], both of which identified key criteria, including usability, reproducibility, scalability, and fault tolerance as essential for assessing workflow system effectiveness.
Di Tommaso et al. [8] specifically addressed reliability in the context of Nextflow, showcasing how the use of containers and reproducible environments can support dependable execution, although debugging and performance tuning remain challenges. The design and adoption of the Common Workflow Language (CWL) are examined by Amstutz et al. [40] and Kotliar et al. [50], who emphasized CWL’s role in standardizing workflow portability across diverse computing environments. Jain et al. [51] introduced FireWorks [51] as a workflow platform tailored for high-throughput computational tasks, addressing the need for dynamic scheduling and job prioritization. Ahmad et al. [52] investigated workflow scheduling challenges, noting that optimizing resource allocation and task dependencies remains a complex problem in large-scale SWS deployments.
Hundreds of SWSs have been developed to support the design, execution, and management of computational workflows across scientific domains [53]. A substantial body of research has examined the core challenges in building and using these systems effectively, including issues of provenance tracking, usability, scalability, and reproducibility [54, 55, 56, 57, 20]. Recent studies have also begun exploring the integration of Large Language Models (LLMs) with SWSs [58, 59, 60], identifying the potential of LLMs, such as GPT models to automate workflow construction, streamline data manipulation, and foster more intelligent, collaborative environments for scientific research. These efforts highlighted the growing convergence between AI and workflow technologies, suggesting promising directions for scalable and user-friendly scientific workflow development.
While prior research has extensively investigated various aspects of SWSs such as reproducibility challenges [44], usability and collaborative development [46, 20] related issues, and workflow portability and reusability limitations [45, 51], these studies have primarily relied on expert reviews, case studies, or controlled experiments. Though such approaches offer valuable system-level insights, they may not fully reflect the day-to-day challenges developers face in practice when designing, maintaining, and adapting SWSs.
To address this gap, our study systematically mined large-scale, real-world developer discussions from GitHub and Stack Overflow to capture practical, community-driven perspectives on SWS development. Notably, we identified topics such as Workflow Execution and Data Structures and Operations as particularly difficult, reflecting prior concerns around reproducibility and data management[29, 43], but we showed new depth by revealing how these difficulties are operationalized and perceived by developers. In addition, we surfaced procedural pain points, such as Distributed Task Management, System Redesign and Managing Rules and Inputs, which remain underrepresented in existing system-level literature but are essential for the operational success of SWSs deployment and maintenance. Furthermore, unlike previous studies that focus on individual systems, our analysis spans a diverse range of SWSs, enabling cross-cutting insights into common bottlenecks and recurring developer questions. This broader scope helps explain observed discrepancies with earlier findings and allows us to offer a more comprehensive understanding of developer needs. By identifying and analyzing recurring pain points through topic modeling, we contributed a complementary, developer-centric perspective for designing more robust, usable, and adaptable workflow systems, currently underrepresented in the literature.